OVHCloud Web Hosting Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#7795 —Optical loop Paris little ring
Scheduled Maintenance Report for Web Cloud
Completed
We have a problem on a fiber. The securisation of some NRA
is no longer active.

For us the problem is between CPY94 and GRA94.

An intervention is being processed.

Update(s):

Date: 2012-12-21 00:42:21 UTC
The team has finished the melting. Everything is secured from our side.


Date: 2012-12-20 23:05:14 UTC
The constraint has been found fixed for OVH, the buckle is again secured. We have lost 1.1 dB thereafter the melting, but this does not impact the good performance of the buckle.

The team on site are melting the fibers of other customers. We are waiting the confirmation of the operator before the closure of the incident.



Date: 2012-12-20 23:02:13 UTC
The constraint on the fiber has not been found yet.
Many difficulties on the field (Clogged or tarred views)
are making the progression so slow.

In new updates, the constraint will be on a portion
of fiber and not on a new connection.
Posted Dec 19, 2012 - 14:27 UTC